# Reinforcement Learning on AYA Dyads to Enhance Medication Adherence

### v1

@@ Appendix A Related Work
Multi-agent RL (MARL). Our proposed approach falls into the range of the independent learners in the MARL literature [16]. Previous literature on MARL in a collaborative game focuses on finding the (approximate) Nash equilibrium of the game through interacting with an unknown environment [34, 8]. However, in our paper, we emphasize the advantage of MARL in terms of its strong interpretability and being able to make decisions in multiple time-scales.
Appendix B Algorithm Details
@@
B.1 Single Agent Algorithm
@@ B.2 MultiAgent Algorithm
Appendix C The Dyadic Environment
C.1 Overview of the Simulated Dyadic Environment
@@
C.1.2 Constructing the dyadic environment.
We follow the environment design in [11], which also uses the Roadmap 2.0 data, but primarily focuses on AYA intervention and relationship intervention. We extend the environment to include the care partner intervention. Specifically, we fit a separate multi-variate linear model for each participant in the dataset with the AR(1) working correlation using the generalized estimating equation (GEE) approach [37, 7]. We impute the treatment effects of the intervention actions based on the typical STE around 0.15, 0.3, and 0.5, which completes a generative model for the state transitions. The environment simulates a trial by randomly sampling dyads from the dataset, and simulate their trajectories based on the actions selected by the RL algorithm. The environment details are described in Appendix C. Our experiments primarily focus on the three vanilla testbeds corresponding to the three STEs.
C.2 Using the Roadmap 2.0 Dataset
@@
To address the missing reward signals, we use directly measurable variables in Roadmap 2.0 dataset as proxies to the outcomes we will observe in the real clinical trial. We approximate AYA adherence, $R_{w,d,t}^{\tiny\texttt{AYA}}$, using the 12-hourly step count from Roadmap 2.0. Previous work has found the two values to be strongly correlated [Hinal:TODO cite]. Since adherence is a binary signal in the ADAPTS-HCT trial, we discretize step count into a binary variable. Furthermore, we approximate the carepartner’s daily psychological distress, $Y_{d}^{\tiny\texttt{CARE}}$, using the daily length of their sleep. Finally, the weekly relationship between the AYA and their carepartner is estimated using the self-reported mood as a surrogate. Specifically, we let $Y_{w}^{\tiny\texttt{REL}}=\mathbbm{1}\{\sum_{d=1}^{7}{\texttt{Mood}}_{w,d}^{\tiny\texttt{AYA}}\geq{\texttt{Mood}}^{\tiny\texttt{AYA}}\}\mathbbm{1}\{\sum_{d=1}^{7}{\texttt{Mood}}_{w,d}^{\tiny\texttt{CARE}}\geq{\texttt{Mood}}^{\tiny\texttt{CARE}}\}$. Here, ${\texttt{Mood}}_{w,d}^{\tiny\texttt{AYA}}$ is the daily self-reported mood on week $w$ and day $d$, and ${\texttt{Mood}}^{\tiny\texttt{AYA}}$ is the $q$-th quantile of the the weekly summed mood across all AYA observations. We choose the quantile level $q$ such that approximately 50% of the dataset satisfies $Y_{w}^{\tiny\texttt{REL}}=1$.
@@
C.3 Environment Model Design
@@
For all transition models, we fit the baseline parameters – which represent system dynamics under no intervention – for each dyad using its respective dataset and a generalized estimating equation [7] approach. We impute the remaining parameters using domain knowledge. Further detail on the choice of the coefficients is in Appendix C.4.
@@
C.4 Selecting Environment Model Parameters
@@
Figure 3 plots the value of the hyperparameter versus the STE computed using the optimal policy in the environment defined by the hyperparameter. We outline our approximation of the optimal policy in Appendix C.5. By default, we choose an environment with mediator effect = 1. This results in three dyadic environments, which we summarize in Table 8.
@@
C.5 Optimal Policy Approximation
@@
C.6 Evidence of the Need for Collaboration in the Dyadic Environment
@@
Appendix D Additional Results
D.1 Ablation Study
@@
D.2 Collaboration of Multi-Agent RL

### v2

@@ Appendix 0.A Related Work
Multi-agent RL (MARL). Our proposed approach falls into the range of the independent learners in the MARL literature [19]. Previous literature on MARL in a collaborative game focuses on finding the (approximate) Nash equilibrium of the game through interacting with an unknown environment [39, 9]. However, in our paper, we emphasize the advantage of MARL in terms of its strong interpretability and being able to make decisions in multiple time-scales. Multi-agent RL has been commonly used in the cooperative game setting, for example, traffic signal control [11, 22]. Our approach is novel in the use of surrogate reward functions for different agents to mitigate the delayed effect of the certain interventions. Our problem is also distinct from the typical cooperative game setting in its asymmetry in different agents’ roles.
Appendix 0.B Algorithm Details
@@
0.B.1 Single Agent Algorithm
@@ 0.B.2 MultiAgent Algorithm
Appendix 0.C The Dyadic Environment
0.C.1 Overview of the Simulated Dyadic Environment
@@
0.C.1.2 Constructing the dyadic environment.
We follow the environment design in [13], which also uses the Roadmap 2.0 data, but primarily focuses on AYA intervention and relationship intervention. We extend the environment to include the care partner intervention. Specifically, we fit a separate multi-variate linear model for each participant in the dataset with the AR(1) working correlation using the generalized estimating equation (GEE) approach [41, 8]. We impute the treatment effects of the intervention actions based on the typical STE around 0.15, 0.3, and 0.5, which completes a generative model for the state transitions. The environment simulates a trial by randomly sampling dyads from the dataset, and simulate their trajectories based on the actions selected by the RL algorithm. The environment details are described in Appendix 0.C. Our experiments primarily focus on the three vanilla testbeds corresponding to the three STEs.
0.C.2 Using the Roadmap 2.0 Dataset
@@
To address the missing reward signals, we use directly measurable variables in Roadmap 2.0 dataset as proxies to the outcomes we will observe in the real clinical trial. We approximate AYA adherence, $R_{w,d,t}^{\tiny\texttt{AYA}}$, using the 12-hourly step count from Roadmap 2.0. Previous work has found the two values to be strongly correlated. Since adherence is a binary signal in the ADAPTS-HCT trial, we discretize step count into a binary variable. Furthermore, we approximate the carepartner’s daily psychological distress, $Y_{d}^{\tiny\texttt{CARE}}$, using the daily length of their sleep. Finally, the weekly relationship between the AYA and their carepartner is estimated using the self-reported mood as a surrogate. Specifically, we let $Y_{w}^{\tiny\texttt{REL}}=\mathbbm{1}\{\sum_{d=1}^{7}{\texttt{Mood}}_{w,d}^{\tiny\texttt{AYA}}\geq{\texttt{Mood}}^{\tiny\texttt{AYA}}\}\mathbbm{1}\{\sum_{d=1}^{7}{\texttt{Mood}}_{w,d}^{\tiny\texttt{CARE}}\geq{\texttt{Mood}}^{\tiny\texttt{CARE}}\}$. Here, ${\texttt{Mood}}_{w,d}^{\tiny\texttt{AYA}}$ is the daily self-reported mood on week $w$ and day $d$, and ${\texttt{Mood}}^{\tiny\texttt{AYA}}$ is the $q$-th quantile of the the weekly summed mood across all AYA observations. We choose the quantile level $q$ such that approximately 50% of the dataset satisfies $Y_{w}^{\tiny\texttt{REL}}=1$.
@@
0.C.3 Environment Model Design
@@
For all transition models, we fit the baseline parameters – which represent system dynamics under no intervention – for each dyad using its respective dataset and a generalized estimating equation [8] approach. We impute the remaining parameters using domain knowledge. Further detail on the choice of the coefficients is in Appendix 0.C.4.
@@
0.C.4 Selecting Environment Model Parameters
@@
Figure 3 plots the value of the hyperparameter versus the STE computed using the optimal policy in the environment defined by the hyperparameter. We outline our approximation of the optimal policy in Appendix 0.C.5. By default, we choose an environment with mediator effect = 1. This results in three dyadic environments, which we summarize in Table 8.
@@
0.C.5 Optimal Policy Approximation
@@
0.C.6 Evidence of the Need for Collaboration in the Dyadic Environment
@@
Appendix 0.D Additional Results
0.D.1 Ablation Study
@@
0.D.2 Collaboration of Multi-Agent RL